Assume the profit earned by an artist in any given year is governed by the function: P(n) = -2n2 + 20n – 30 where n represents the number of paintings sold and P (in thousands of dollars) represents the profit. How many paintings should be sold to create a maximum profit? What is the maximum profit?

Answer by stanbon(75887)

Assume the profit earned by an artist in any given year is governed by the function: P(n) = -2n2 + 20n – 30 where n represents the number of paintings sold and P (in thousands of dollars) represents the profit. How many paintings should be sold to create a maximum profit? What is the maximum profit?
------
P is a quadratic with a = -2 and b = 20 (number of paintings need to be sold)
Maximum P occurs at n = -b/2a = -20/-4 = 5
--------------------------
P(5) = -2(5)^2 + 20*5 - 30 = -50 + 100 - 30
P(5) = $20 (maximum profit)
==============================
Cheers,
Stan H.